OpenStudy (anonymous):

find the real or imaginary solutions by factoring? x^4 - 3x^2 = -2x^2

OpenStudy (anonymous):

add 2x^2 to both sides what you get??

OpenStudy (anonymous):

hi there please re-state the problem i'm having problems analyzing what your really asking for?

OpenStudy (anonymous):

find the real or imaginary solutions by factoring? x^4 - 3x^2 = -2x^2

OpenStudy (anonymous):

that's all it says

OpenStudy (anonymous):

now add 2x^2 to both sides what you get???

OpenStudy (anonymous):

its easy do it

OpenStudy (anonymous):

thats not how you answer these questions. I will help you to get the correct answer. First we will use the factor theorem. Have you ever learnt about the factor theorem?

OpenStudy (anonymous):

ok lets do it

OpenStudy (anonymous):

you have this x^4 - 3x^2 = -2x^2 adding 2x^2 to both sides we get x^4-x^2=0 x^2(x^2-1)=0 now either x^2 = 0 or x^2-1 = 0 can you find the solution now???

OpenStudy (anonymous):

x= -1. Plug in -1 for x in the equation and the equation will be equal to zero.

OpenStudy (anonymous):

if x^2 = 0 then x.x = 0 x=0 , x =0 if x^2 - 1 = 0 then x^2 = 1 x=+1 , x = -1 thus all the solutions are real the possible solutions are 0,0,1,-1